OpenStudy (anonymous):

hamburgers for $1.50 and hot dogs sell for $1.25 if 78 hot dogs and hamburgers were sold and brought 105.50 what is the exact number of hotdogs

OpenStudy (anonymous):

Do u still need help??? I would be glad to...

OpenStudy (anonymous):

Well even if u got it. Here it goes... let x=number of hamburgers And.. 78+x=number of hotdogs Hamburgers are $1.50 and hot dogs are $1.25 so... 1.50(x) +1.25(78+x)=105.50 1.50x +97.5 +1.25x=105.50 2.75x=105.50-97.5 2.75x=8 2.75x=8 2.75x 2.75 x=3 Therefore 3=number of hamburgers and 81=number of hotdogs.:) There you have it! Please give metal if I was of any help.:)Thanks and good luck.